# Automatic Depression Assessment using Machine Learning: A Comprehensive Survey

### v1

@@ 4.3. Face/head-based depression assessment
Due to ethical considerations, some ADA studies are developed to infer depression from anonymous facial primitives. For instance, Pan et al. (Pan et al., 2023) apply CNN to learn a video-level representation from facial landmarks, and then propose a spatio-temporal attention network to extract its depression-related features for ADA. Niu et al. (Niu et al., 2024a) first divide the video-level facial landmark sequence into several segments along the temporal dimension, and then extend the MLP-mixer (Tolstikhin et al., 2021) to make a depression prediction from every segment. Kacem et al. (Kacem et al., 2018) utilise GMM model and Fisher vector encoding to represent Lie-algebra based rotation matrix of 3D head motions. Based on eye blinks, Al et al. (Al-gawwam and Benaissa, 2018) summarise video-level eye blink-related features such as blink rate and duration from facial landmarks, and then individually employ SVM, Naive Bayes and Adaboost for binary depression classification. Since a systematic study investigating the relationship between AUs and depression (Cohn et al., 2009) suggests that AUs can provide vital information for depression assessment, a certain number of ADA systems (Song et al., 2022a, 2018; Gong and Poellabauer, 2017) also attempt to extract informative depression cues from AUs, e.g., Muzammel et al. (Muzammel et al., 2021) utilise a MLP to process 20 AUs, while Gong et al. (Gong and Poellabauer, 2017) computing statistics of AUs appeared in every interview topic for for video-level ADA.
@@
Instead of using only a single primitive, more approaches utilise multiple facial primitives for ADA. The baselines of AVEC 2016 and AVEC 2017 (Valstar et al., 2016; Ringeval et al., 2017) feed various frame-level facial primitives (e.g., AUs, head poses, gaze directions and discrete emotions) to linear SVM or RF for frame-level depression classification orseverity estimation. Then, the video-level depression classification or severity estimation is achieved via temporal majority voting, averaging frame-level predictions, or aggregating facial primitive statistics. Song et al. (Song et al., 2018) also compute 12 statistics to summarise all frame-level AUs/head poses/gaze directions of the given video, which are concatenated and processed by SVR for video-level ADA. Guo et al. (Guo et al., 2022) propose a temporal dilated CNN to extract long-term video-level facial dynamics from frame-level facial landmark and head pose sequences, and then apply an attention operation to adaptively weight different feature map channels for video-level ADA. Niu et al. (Niu et al., 2024b) propose MLPs with a Dual Gating to process facial landmarks and AUs of the given video for depression severity estimation, where the Dual Gating maintains the physical properties of input landmarks and AUs while a Mutual Guidance module interacting facial landmarks and AUs in the context of their global information. Recently, Islam et al. (Islam and Bae, 2024) develop an mobile system for real-world ADA, where segment-level statistics of facial landmarks, AUs, gazes and head poses are fused and selected for inferring depression.
4.4. Body behaviour-based depression assessment
@@
Human body behaviours, especially gait behaviours, are distinguishable between depressed and non-depressed individuals (Wang et al., 2020b, 2021b). Some studies (Wang et al., 2020a, 2021b) identified that silhouette extraction is a key challenge unique to body behaviour-based ADA. To address this, they learn depression-related time-domain features for analysing gait abnormalities, frequency-domain features describing video-level body dynamics and frame-level spatial geometric features, which are finally concatenated and fed to a MLP or SVM for ADA. Besides, Shao et al. (Shao et al., 2021) leverage LSTM to process sequential skeleton behaviours while two CNNs are trained to extract silhouette features from the gait video. The predictions of both modalities are combined to make the final depression prediction. Liu et al. (Liu et al., 2022) develop a two-stream 3D CNN to predict depression severity from human body movements during walking, i.e., the motion amplitude computed from the 18 joints’ positions of the target individual. These positions are fused with deep-learned spatio-temporal facial expression cues for video-level ADA.
@@
In addition to gait, limb movements are also important biomarker for assessing depression. Joshi et al. (Joshi et al., 2012) compute spatio-temporal cues from human upper body key points via a temporal visual words dictionary which are fused with LBP-TOP feature-based facial behaviour codebook learned from the video, where MLP and SVM-based classifiers are employed to make the final depression assessment. Their later work (Joshi et al., 2013a) further explores relative orientation and radius of the human upper body in addition to the upper body key points, which are then fused and processed by SVM for ADA. Yu et al. (Yu et al., 2022) propose a spatial attention dilated TCN (SATCN) which applies a hierarchy group of temporal convolution operations with various dilated convolution scales to learn human skeletal cues for ADA. Lu et al. (Lu et al., 2022) extract a rigid-body representation by analysing kinetic energy and potential energy collected from the human walking video, and then apply fast Fourier transform to encode them as a joint energy feature, where SVM, RF and KNN are individually employed as the classifier. Ahmad et al. (Ahmad et al., 2021) individually apply various 2D CNNs to assess depression from either human upper body frames or facial frames, and then averaging all frame-level predictions to achieve video-level ADA, which suggests that face and upper body frames provide similar ADA performance. Lin et al. (Lin et al., 2021) leverage various body behaviours for ADA by extracting frame-level body pose features using OpenPose. From these, they derive multiple body behaviour features, including video-level statistics, self-adaptor features reflecting individual body manipulation, and fidgeting behaviour features indicating unconscious or semi-conscious movements. A GMM and Fisher Vector encoding are then used to summarise frame-level features into a video-level representation for ADA.
4.5. Audio-based depression assessment
@@
While human non-verbal audio behaviours contain abundant depression-indicative information (Cummins et al., 2015), three types of audio features have been widely explored for ADA, including: (i) acoustic spectral and frequency features (e.g., spectral, frequency, voice quality and glottal features); (ii) prosodic features measuring quasi-periodic rate of vocal fold vibration, speaking rate, pausing and intonation patterns (Moore II et al., 2007); and (iii) deep-learned features capturing human uninterpretable but depression-specific audio cues.
4.5.1. Challenges.
@@
To extract hand-crafted audio features for ADA, (i) feature pre-processing is challenging as these features may possess high dimensionality. Jan et al. (Jan et al., 2017) apply PCA to the extracted MFCC features along with various spectral LLDs (e.g., spectral flatness, spectral flux, and other spectral features). Zhang et al. (Zhang et al., 2024a) vertically concatenate visual representations of complementary MFCCs and mel spectrogram extracted from original audio waveform to form a matrix, which is subsequently processed by a 2D-CNN, an attention mechanism, and a BiLSTM for ADA. Alternatively, DL-based approaches (discussed in Sec. 4.5.3) have been proposed to tackle the (ii) unaddressed problem posed by hand-crafted features, i.e., they can not capture subtle and human uninterpretable depression indicators within audio signals. Besides, (iii) aggregating frame/segment-level audio feature as a clip-level audio depression representation is another key research question (Shen et al., 2017), where statistics (Moore II et al., 2007), GMM (Williamson et al., 2016), temporal convolution (Ye et al., 2021), temporal attention (Pan et al., 2023), LSTM (Zhang et al., 2021; Ma et al., 2016; Al Hanai et al., 2018), majority voting (Vázquez-Romero and Gallardo-Antolín, 2020), and GNNs (Chen et al., 2023) have been extended to address this challenge. We detailed survey the existing audio-based ADA methods within the subsections below.
4.5.2. Hand-crafted audio feature-based approaches:
@@
Mel-frequency cepstral coefficients (MFCCs) (Zhang et al., 2024a; Othmani et al., 2021; He and Cao, 2018; Rejaibi et al., 2022; Jan et al., 2017; Gupta et al., 2014; Fu et al., 2022; Pérez-Toro et al., 2024) is one of the most widely adopted acoustic spectral and frequency features for ADA, which are derived from spectrograms by capturing the short-term energy spectrum and revealing subtle changes in energy overtime, such as monotony, slower speech rate, and increase pauses (related to depression) (Kraepelin, 1921). As a result, a large number of studies (Chen et al., 2023; Yang et al., 2017; Williamson et al., 2016) leveraged MFCCs along with other hand-crafted acoustic low-level descriptors (LLDs) for ADA. The results validate that the MFCCs are the best individual indicator for ADA among different audio spectral and frequency features. Speeches expressed by depressed individuals are typically characterised by lacking in expression and reduced variability in fundamental frequency and prosody (Kraepelin, 1921), voice perceptual qualities, prosodic/vocal tract features (e.g., movement of articulators), pitch (emotional and physiological changes in speeches (Hollien, 1980; Gupta, 2009)) as well as glottal waveform relating to the volume–velocity of airflow through the glottis from the lungs (Ozdas et al., 2004). Consequently, these acoustic features have been specifically employed by audio/multi-modal ADA systems (Zhang et al., 2024a; Muzammel et al., 2020; Sun et al., 2017). Moore et al. (Moore II et al., 2007) extract prosodic, vocal tract, and glottal source features at the frame level to analyse audio recordings from English-speaking participants reading a short story for at least three minutes. Clip-level representations are then obtained using mean, standard deviation, and temporal summaries of these frame-level features. A quadratic discriminant analysis (QDA)-based classifier finally performs depression classification using each audio clip-level representation. Results indicate that the combination of prosodic and glottal features yields the highest accuracy. Williamson et al. (Williamson et al., 2016) utilise a formant tracking algorithm (Mehta et al., 2012) to estimate F1, F2 and F3 from each normalised audio clip in every 10 ms. Meanwhile, they employ peak-to-root-mean-square (RMS) to measure loudness, and openSMILE to extract MFCCs on a segment-level basis. All these features are jointly fed to a Gaussian staircase regressor to learn the entire audio clip-level depression representation. Similarly, Alhanai et al. (Al Hanai et al., 2018) also apply COVAREP to extract frame-level formant, pitch, glottal and voice quality features, which are fed to a three-layer LSTM for audio clip-level ADA.
4.5.3. Deep learning approaches.
@@
DL-based approaches are frequently achieved through three strategies: (i) leveraging/fine-tuning existing pre-trained DL networks; (ii) developing specific DL networks for preliminary depression-related audio feature extraction; and (iii) developing specific DL networks to further refine extracted hand-crafted audio features. In particular, the DeepSpectrum (Amiriparian et al., 2017; Zhao et al., 2019), SincNet (Ravanelli and Bengio, 2018), and NetVLAD (Arandjelovic et al., 2016) are the most widely-used pre-trained DL networks for audio-based ADA, as they provide strong emotion-related audio representations. An typical example is proposed by (Ye et al., 2021), which applies DeepSpectrum to convert raw audio waveform as a set of segment-level spectrograms, and then extract segment-level deep features via a pre-trained VGG-16. The clip-level depression representation is finally aggregated via a Temporal Convolutional Network (TCN) with an attention layer. Pan et al. (Pan et al., 2023) adopt the pre-trained SincNet to reduce the amount of model training needed to extract frame-level audio features. Then, a Temporal Attention (TA) block adaptively weights such frame-level features along the temporal dimension, which are further fed to an unaligned transformer to generate the clip-level audio representation for ADA. Instead of using pre-trained models, Zhang et al. (Zhang et al., 2021) trained a self-supervised DEPA model using the Switchboard dataset (Godfrey et al., 1992). The trained DEPA processes pre-extracted audio Mel-scale spectrograms through CNN layers, followed by an LSTM for clip-level ADA. A similar approach (Vázquez-Romero and Gallardo-Antolín, 2020) also utilises segment-level audio spectrograms as the input to a 1D-CNN, where the majority voting is applied to aggregate segment-level predictions for clip-level depression classification to address the clip-level aggregation challenge. Othmani et al. (Othmani et al., 2021) propose an EmoAudioNet which starts with extracting frame-level MFCCs and segment-level spectrogram features, and then individually processes them via a MFCC-based CNN and a spectrogram-based CNN, whose outputs are finally concatenated for depression assessment. Chen et al. (Chen et al., 2023) split each audio signal into a set of one-second audio segments, from each of which MFCCs and other hand-crafted LLD features are extracted. These segment-level audio features are subsequently processed by a GNN-based Semi-Supervised Domain Adaptation model (GNN-SDA) for ADA. In (Ma et al., 2016), each extracted frame-level Mel-scale filter bank feature first undergoes a standard normal variate normalization, and are then concatenated along time-axis as a set of segment-level 2D time-frequency representations, which are further processed via a CNN-LSTM model and a final SGD regressor to make segment-level depression prediction.
4.5.4. Hybrid approaches.
@@
To leverage advantages of both hand-crafted and deep-learned features for a more effective depression assessment, Kaya et al. (Kaya et al., 2019) utilise the extended Geneva minimalistic acoustic parameter set (eGeMAPs) (Eyben et al., 2015) to extract frame-level LLDs (e.g., breathing, fillers, lip noise, laughter) and apply DeepSpectrum to learn segment-level audio features. These feature sets are then combined for ADA, with results highlighting the complementarity of hand-crafted and deep-learned audio features in ADA. He et al. (He and Cao, 2018) extract 38 hand-crafted LLDs capturing spectral, cepstral, prosodic and voice quality cues, while deep learning (via a CNN) features from spectrograms and waveforms of the given audio clip. Based on them, semantic median robust extended local binary patterns (MRELBP) are further computed. All features are concatenated and passed to a FC layer for ADA. Yang et al. (Yang et al., 2017) extract hand-crafted frame-level MFCCs, loudness, energy, pitch, jitter and shimmer features from the given audio, and further summarise them as a clip-level statistical representations. Finally, a CNN is employed to learn the clip-level depression score. Rejaibi et al. (Rejaibi et al., 2022) extract MFCCs, 53 hand-crafted acoustic features (e.g., fundamental frequency (F0), formants, jitter and shimmer). These frame-level features are then fed to an RNN model to learn their temporal dependencies and produce the clip-level depression prediction.
4.6. Language/text-based depression assessment

### v2

@@ 4.3.4 Facial behaviour primitive-based approaches
Due to ethical concerns, original face images/videos sometimes are not available, as faces reveal identity while depression status is sensitive. Thus, some ADA studies are developed to infer depression from anonymous facial attributes, where frame-level facial landmarks, discrete facial expressions, Action Units (AUs), head poses/motions, and eye gaze directions and blinks are the most prevalent primitives. For instance, Pan et al. [35] apply CNN to learn a video-level representation from facial landmarks, and then propose a spatio-temporal attention network to extract its depression-related features for ADA. Niu et al. [197] first divide the video-level facial landmark sequence into several segments along the temporal dimension, and then extend the MLP-mixer [281] to make a depression prediction from every segment. Kacem et al. [282] utilise GMM model and Fisher vector encoding to represent Lie-algebra based rotation matrix of 3D head motions. Based on eye blinks, Al et al. [174] summarise video-level eye blink-related features such as blink rate and duration from facial landmarks, and then individually employ SVM, Naive Bayes and Adaboost for binary depression classification. Since a systematic study investigating the relationship between AUs and depression [117] suggests that AUs can provide vital information for depression assessment, a certain number of ADA systems [42, 214, 36] also attempt to extract informative depression cues from AUs, e.g., Muzammel et al. [26] utilise a MLP to process 20 AUs, while Gong et al. [36] computing statistics of AUs appeared in every interview topic for for video-level ADA.
@@
Instead of using only a single primitive, more approaches utilise multiple facial primitives for ADA. The baselines of AVEC 2016 and AVEC 2017 [61, 62] feed various frame-level facial primitives (e.g., AUs, head poses, gaze directions and discrete emotions) to linear SVM or RF for frame-level depression classification orseverity estimation. Then, the video-level depression classification or severity estimation is achieved via temporal majority voting, averaging frame-level predictions, or aggregating facial primitive statistics. Song et al. [214] also compute 12 statistics to summarise all frame-level AUs/head poses/gaze directions of the given video, which are concatenated and processed by SVR for video-level ADA. Guo et al. [185] propose a temporal dilated CNN to extract long-term video-level facial dynamics from frame-level facial landmark and head pose sequences, and then apply an attention operation to adaptively weight different feature map channels for video-level ADA. Niu et al. [186] propose MLPs with a Dual Gating to process facial landmarks and AUs of the given video for depression severity estimation, where the Dual Gating maintains the physical properties of input landmarks and AUs while a Mutual Guidance module interacting facial landmarks and AUs in the context of their global information. Recently, Islam et al. [215] develop an mobile system for real-world ADA, where segment-level statistics of facial landmarks, AUs, gazes and head poses are fused and selected for inferring depression.
4.4 Body behaviour-based depression assessment
@@
Human body behaviours, especially gait behaviours, are distinguishable between depressed and non-depressed individuals [283, 178]. Some studies [177, 178] identified that silhouette extraction is a key challenge unique to body behaviour-based ADA. To address this, they learn depression-related time-domain features for analysing gait abnormalities, frequency-domain features describing video-level body dynamics and frame-level spatial geometric features, which are finally concatenated and fed to a MLP or SVM for ADA. Besides, Shao et al. [188] leverage LSTM to process sequential skeleton behaviours while two CNNs are trained to extract silhouette features from the gait video. The predictions of both modalities are combined to make the final depression prediction. Liu et al. [198] develop a two-stream 3D CNN to predict depression severity from human body movements during walking, i.e., the motion amplitude computed from the 18 joints’ positions of the target individual. These positions are fused with deep-learned spatio-temporal facial expression cues for video-level ADA.
@@
In addition to gait, limb movements are also important biomarker for assessing depression. Joshi et al. [179] compute spatio-temporal cues from human upper body key points via a temporal visual words dictionary which are fused with LBP-TOP feature-based facial behaviour codebook learned from the video, where MLP and SVM-based classifiers are employed to make the final depression assessment. Their later work [180] further explores relative orientation and radius of the human upper body in addition to the upper body key points, which are then fused and processed by SVM for ADA. Yu et al. [189] propose a spatial attention dilated TCN (SATCN) which applies a hierarchy group of temporal convolution operations with various dilated convolution scales to learn human skeletal cues for ADA. Lu et al. [181] extract a rigid-body representation by analysing kinetic energy and potential energy collected from the human walking video, and then apply fast Fourier transform to encode them as a joint energy feature, where SVM, RF and KNN are individually employed as the classifier. Ahmad et al. [40] individually apply various 2D CNNs to assess depression from either human upper body frames or facial frames, and then averaging all frame-level predictions to achieve video-level ADA, which suggests that face and upper body frames provide similar ADA performance. Lin et al. [67] leverage various body behaviours for ADA by extracting frame-level body pose features using OpenPose. From these, they derive multiple body behaviour features, including video-level statistics, self-adaptor features reflecting individual body manipulation, and fidgeting behaviour features indicating unconscious or semi-conscious movements. A GMM and Fisher Vector encoding are then used to summarise frame-level features into a video-level representation for ADA.
4.5 Audio-based depression assessment
@@
While human non-verbal audio behaviours contain abundant depression-indicative information [136], three types of audio features have been widely explored for ADA, including: (i) acoustic spectral and frequency features (e.g., spectral, frequency, voice quality and glottal features); (ii) prosodic features measuring quasi-periodic rate of vocal fold vibration, speaking rate, pausing and intonation patterns [216]; and (iii) deep-learned features capturing human uninterpretable but depression-specific audio cues.
4.5.1 Challenges and solutions
@@
To extract hand-crafted audio features for ADA, (i) feature pre-processing is challenging as these features may possess high dimensionality. Jan et al. [265] apply PCA to the extracted MFCC features along with various spectral LLDs (e.g., spectral flatness, spectral flux, and other spectral features). Zhang et al. [270] vertically concatenate visual representations of complementary MFCCs and mel spectrogram extracted from original audio waveform to form a matrix, which is subsequently processed by a 2D-CNN, an attention mechanism, and a BiLSTM for ADA. Alternatively, DL-based approaches (discussed in Sec. 4.5.3) have been proposed to tackle the (ii) unaddressed problem posed by hand-crafted features, i.e., they can not capture subtle and human uninterpretable depression indicators within audio signals. Besides, (iii) aggregating frame/segment-level audio feature as a clip-level audio depression representation is another key research question [233], where statistics [216], GMM [268], temporal convolution [243], temporal attention [35], LSTM [220, 219, 224], majority voting [218], and GNNs [271] have been extended to address this challenge. We detailed survey the existing audio-based ADA methods within the subsections below.
4.5.2 Hand-crafted audio feature-based approaches
@@
Mel-frequency cepstral coefficients (MFCCs) [270, 229, 223, 230, 265, 251, 228, 236] is one of the most widely adopted acoustic spectral and frequency features for ADA, which are derived from spectrograms by capturing the short-term energy spectrum and revealing subtle changes in energy overtime, such as monotony, slower speech rate, and increase pauses (related to depression) [126]. As a result, a large number of studies [271, 274, 268] leveraged MFCCs along with other hand-crafted acoustic low-level descriptors (LLDs) for ADA. The results validate that the MFCCs are the best individual indicator for ADA among different audio spectral and frequency features. Speeches expressed by depressed individuals are typically characterised by lacking in expression and reduced variability in fundamental frequency and prosody [126], voice perceptual qualities, prosodic/vocal tract features (e.g., movement of articulators), pitch (emotional and physiological changes in speeches [23, 17]) as well as glottal waveform relating to the volume–velocity of airflow through the glottis from the lungs [217]. Consequently, these acoustic features have been specifically employed by audio/multi-modal ADA systems [270, 231, 222]. Moore et al. [216] extract prosodic, vocal tract, and glottal source features at the frame level to analyse audio recordings from English-speaking participants reading a short story for at least three minutes. Clip-level representations are then obtained using mean, standard deviation, and temporal summaries of these frame-level features. A quadratic discriminant analysis (QDA)-based classifier finally performs depression classification using each audio clip-level representation. Results indicate that the combination of prosodic and glottal features yields the highest accuracy. Williamson et al. [268] utilise a formant tracking algorithm [284] to estimate F1, F2 and F3 from each normalised audio clip in every 10 ms. Meanwhile, they employ peak-to-root-mean-square (RMS) to measure loudness, and openSMILE to extract MFCCs on a segment-level basis. All these features are jointly fed to a Gaussian staircase regressor to learn the entire audio clip-level depression representation. Similarly, Alhanai et al. [224] also apply COVAREP to extract frame-level formant, pitch, glottal and voice quality features, which are fed to a three-layer LSTM for audio clip-level ADA.
4.5.3 Deep learning approaches
@@
DL-based approaches are frequently achieved through three strategies: (i) leveraging/fine-tuning existing pre-trained DL networks; (ii) developing specific DL networks for preliminary depression-related audio feature extraction; and (iii) developing specific DL networks to further refine extracted hand-crafted audio features. In particular, the DeepSpectrum [225, 226], SincNet [285], and NetVLAD [286] are the most widely-used pre-trained DL networks for audio-based ADA, as they provide strong emotion-related audio representations. An typical example is proposed by [243], which applies DeepSpectrum to convert raw audio waveform as a set of segment-level spectrograms, and then extract segment-level deep features via a pre-trained VGG-16. The clip-level depression representation is finally aggregated via a Temporal Convolutional Network (TCN) with an attention layer. Pan et al. [35] adopt the pre-trained SincNet to reduce the amount of model training needed to extract frame-level audio features. Then, a Temporal Attention (TA) block adaptively weights such frame-level features along the temporal dimension, which are further fed to an unaligned transformer to generate the clip-level audio representation for ADA. Instead of using pre-trained models, Zhang et al. [220] trained a self-supervised DEPA model using the Switchboard dataset [287]. The trained DEPA processes pre-extracted audio Mel-scale spectrograms through CNN layers, followed by an LSTM for clip-level ADA. A similar approach [218] also utilises segment-level audio spectrograms as the input to a 1D-CNN, where the majority voting is applied to aggregate segment-level predictions for clip-level depression classification to address the clip-level aggregation challenge. Othmani et al. [229] propose an EmoAudioNet which starts with extracting frame-level MFCCs and segment-level spectrogram features, and then individually processes them via a MFCC-based CNN and a spectrogram-based CNN, whose outputs are finally concatenated for depression assessment. Chen et al. [271] split each audio signal into a set of one-second segments, from each of which MFCCs and other hand-crafted LLD features are extracted. These segment-level audio features are subsequently processed by a GNN-based Semi-Supervised Domain Adaptation model (GNN-SDA) for ADA. In [219], each extracted frame-level Mel-scale filter bank feature first undergoes a standard normal variate normalization, and are then concatenated along time-axis as a set of segment-level 2D time-frequency representations, which are further processed by a CNN-LSTM model and a SGD regressor to make segment-level depression predictions.
4.5.4 Hybrid approaches.
@@
To leverage advantages of both hand-crafted and deep-learned features for a more effective audio-based ADA, Kaya et al. [256] utilise the extended Geneva minimalistic acoustic parameter set (eGeMAPs) [288] to extract frame-level LLDs (e.g., breathing, fillers, lip noise, laughter) and apply DeepSpectrum to learn segment-level audio features. These feature sets are then combined for ADA, with results highlighting the complementarity of hand-crafted and deep-learned audio features in ADA. He et al. [223] extract 38 hand-crafted LLDs capturing spectral, cepstral, prosodic and voice quality cues, while deep learning (via a CNN) features from spectrograms and waveforms of the given audio clip. Based on them, semantic median robust extended local binary patterns (MRELBP) are further computed. All features are concatenated and passed to a FC layer for ADA. Yang et al. [274] extract hand-crafted frame-level MFCCs, loudness, energy, pitch, jitter and shimmer features from the given audio, and further summarise them as a clip-level statistical representations. Finally, a CNN is employed to learn the clip-level depression score. Rejaibi et al. [230] extract MFCCs, 53 hand-crafted acoustic features (e.g., fundamental frequency (F0), formants, jitter and shimmer). These frame-level features are then fed to an RNN model to learn their temporal dependencies and produce the clip-level depression prediction.
4.6 Language/text-based depression assessment